Temple Inc. has equipment with a carrying amount of $800,000. The equipment’s fair value less costs...

  1. Temple Inc. has equipment with a carrying amount of $800,000. The equipment’s fair value less costs to sell is $780,000, and its value-in-use is $815,000. The equipment is expected to be used in operations in the future. What amount should Temple Inc. report as an impairment to its equipment? $
  1. Ludo Inc. acquires a patent for a drug with a remaining legal and useful life of six years on January 1, 2014 for $2,100,000. The company uses straight-line amortization for patents. On January 2, 2016, a new patent is received for a timed-release version of the same drug. The new patent has a legal and useful life of twenty years. The least amount of amortization that could be recorded in 2016 is $ Answer
  1. Ricco Co. purchased machinery on January 2, 2010, for $440,000. The straight-line method is used and useful life is estimated to be 10 years, with a $40,000 residual value. At the beginning of 2016 Ricco spent $96,000 to overhaul the machinery. After the overhaul, Ricco estimated that the useful life would be extended 4 years (14 years total), and the residual value would be $20,000. The depreciation expense for 2016 should be $ Answer
Answer #1

1. Carrying value of the equipment is given = $ 800,000. The impairment of the assets arises when the carrying value is higher than of its fair value less cost sto sell or value in use.

On the given case fair value less cost to sell is $ 780,000 bit value of future use = $ 815,000 so value to be taken for impairment analysis will be higher of this two i.e value in use of $ 815,00/.

Now, Value in use is higher than the carrying value so there is no need for the impairment of assets.

Temple Inc. report as an impairment to its equipment - $ 0

2. The patents are intangible assets to be ammortized over its useful year life on a straight line basis, so for the assets acquired on 1st Jan 2014 for 2,100,000 will be ammortized accordingly i.e $ 2,100,000 / 6 = $ 350,000.

So value of patent on 2nd jan 2016. Will be cost - ammortization i.e ( $ 2100000 - 350000-350000) = $ 1,400,000. Which will be ammortized over the revised useful life.

Here the new patent value in use is 20 years , so the remaining value of $ 1,400,000 will be ammortized over the 20 years i.e $ 70,000 ( 1400000 / 20) every year.

Thus The least amount of amortization that could be recorded in 2016 is $ 70,000.

3.

Rocco. Co assets cost = $ 440,000 useful life = 10 years and residual value is = $ 40,000.

Thus straight line depreciation every year =( Cost - salvage value) / useful life.

Straight line depreciation every year = ( $ 440,000 - $ 40,000 ) / 10. = 400000/ 10

Straight line depreciation every year = $ 40,000.

Thus depreciation till 1st Jan 2026 will be = 40,000 × 6 = $ 240,000.

Remaining value as on 1st Jan 2016 = $ 200,000 ( 440000 - 240000,).

Overhaul Expenses when increase the useful life then it is eligible for capitalisation of the assets i.e added to the value of assets.

Thus value of assets as on 2nd jan 2016 = $ 200,000 + 96,000 = $ 296,000.

Depreciation for 2016 = ( $ 296,000 - $ 20,000) / 8.

Depreciation for 2016 = $ 276,000 /8 = $ 34,500.

Thus, The depreciation expense for 2016 should be $ 34,500.

Feel free to comment doubts if any..